# Bispectrum Islands

### v1

@@ VI Examples of explicit scenarios
Our bounds must be satisfied by any bispectrum resulting from the tree-level exchange of a unitary sector coupled to the inflaton field as specified in (11). The simplest scenario to test involves a linear superposition of a finite number of resonances in the hidden sector. In Figure 6, we illustrate a few examples with three such resonances and confirm that the resulting bispectrum shapes lie within the allowed region. We emphasize again that, as shown by the red curve in the figure, the bispectrum can indeed change sign for sufficiently squeezed triangle configurations. This behavior arises from the cosmological collider oscillations dominating over the local contributions to the bispectrum (see Eq. (31)).
@@
As another non-trivial example, we take the hidden sector to be a conformal field theory defined in the bulk of de Sitter space, with ${\cal O}$ representing a scalar CFT operator of the scaling dimension $\delta$. (Such CFT should not be confused with a hypothetical CFT defined on the future boundary of de Sitter space, which might be dual to the theory in the bulk Strominger (2001).)
@@
×$\times\Gamma(\delta-3/2-i\mu)\,\Gamma(\delta-3/2+i\mu)\,.$ (53)
@@
In the case of a CFT operator with conformal dimension $\delta$, the spectral density asymptotically behaves as $\rho(\mu)\propto\mu^{2\delta-3}$. This implies that, for the convergence of the spectral integral, we must have $\delta<2$. In conclusion, we expect our bounds to hold for the exchange of scalar operators with dimensions in the range 3/$23/2<\delta<$ 2, where the two-point function can be expressed as a UV-finite integral over the mass spectrum of heavy fields in principal series.
@@
In Figure 7, using the spectral density given in Eq. (53), we have numerically evaluated the bispectrum arising from the exchange of a scalar CFT operator for three sample values of the conformal dimension $\delta\in(3/2,2)$. As shown in the plots, the resulting shapes lie within the allowed region.
@@ VIII Typical non-Gaussianity templates and beyond
It is worth noting that although the equilateral and orthogonal templates themselves are outside the allowed region, they still provide a useful basis for effectively capturing many of the curves within the peninsula (corresponding to different choices of the spectral density $\rho(\mu)$). As a result, existing observational bounds on $f_{\text{NL}}^{\text{eq}}$ and $f_{\text{NL}}^{\text{orth}}$ can already be used to constrain the amplitude of such curves, along the lines of Cabass et al. (2024).
Looking ahead, it would be interesting to identify a sufficient set of templates to effectively capture the diversity of possible shapes encompassed by the peninsula (in this direction, see the recent work Sohn et al. (2024) where a variety of templates associated with the exchange of massive fields are compared with the Planck CMB data—though some of the studied examples lie outside the scope of our assumptions). Beyond the standard template-based analyses, it is also important to consider how best to analyze the bispectrum data to assess whether it favors or disfavors the carved region as a whole. We leave this question for future investigations.
@@ IX Future Directions
Figure 8: The figure shows that the standard non-Gaussianity templates (Local, Orthogonal, and Equilateral) are inconsistent with our bounds, while the shape generated by the leading-order EFT operator $\delta\dot{\phi}(\partial\delta\phi)^{2}$ lies within our bispectrum island. See the discussions in Section IX.
@@
Acknowledgments We would like to thank Giovanni Cabass, Sebastián Céspedes, Calvin Yi-Ren Chen, Carlo Contaldi, Enrico Pajer, Guillermo Pimentel, Arthur Poisson, David Stefanyszyn, Sébastien Renaux-Petel, Xi Tong, Kamran Salehi Vaziri, Don Gang Wang, Denis Werth,Chen Yang and Yuhang Zhu for stimulating discussions. We also thank Borna Salehian for comments on the manuscript. CdR and AJT are partially supported by the STFC Consolidated Grant ST/X000575/1. CdR and SJ are also supported by Simons Investigator Award 690508.
@@ Appendix A Large Boost Breaking Interactions
Figure 10: The left panel shows that in the Type II setup studied in Appendix A, the bispectrum template $B(k_{i};\mu)$ evaluated at the equilateral configuration changes sign as $\mu$ varies. In contrast, for certain triangle shapes—such as $(1,1,0.1)$—the sign remains fixed for all values of $\mu$ (right). See Appendix A for further discussion.
@@
$M_{\mathrm{Pl}}\to\infty\,,\,\,\dot{H}\to 0\,\,\,(M_{\mathrm{Pl}}^{2}|\dot{H}|=\text{const.})\,,$. ) , (63)
@@
Notably, unlike the Type I setup, in Type II we cannot establish any definite sign for the bispectrum based on the positivity of $\rho(\mu)$. The reason is that the overall coupling $\kappa_{*}$ appearing above is not sign definite. In fact, even with a definite sign for $\kappa_{*}$, it is not possible to put bounds on the sign of the equilateral bispectrum in Type II, because the template $B(1,1,1;\mu)$ changes sign as $\mu$ is varied, see Figure 10. This is also a roadblock to putting bounds on the bispectrum shape, conventionally normalized at the equilateral configuration. Such object would be unbounded for the same reason.
@@
which is expected to be bounded from above and below. Akin to the bispectrum shape in Type I, at a fixed momentum configuration, the shape $\tilde{{\cal S}}(k_{i})$ is also optimized by a spectral density that sharply picks at a particular mass-index $\mu=\mu_{\text{max,min}}$, which we find numerically. In other words, we find the optimal values of the ansatz
@@
by varying the mass across the entire principal series, i.e. $\mu\in[0,\infty)$. As shown in Figure 9, unlike the bounds obtained in the Type I setup, the maximum value of $\tilde{{\cal S}}$ is not set by the bispectrum associated with the exchange of the lightest field (i.e. $\mu=0$). Instead, for configurations with $k_{3}/k_{1}\gtrsim 0.1$, we find that $\mu_{\text{max}}\approx 1.9$ while $\mu_{\text{min}}=0$. For more squeezed triangles, however, $\mu_{\text{max}}$ decreases to zero, while $\mu_{\text{min}}$ varies from 0 in the squeezed limit up to $\approx 1.6$ as $k_{3}/k_{1}$ approaches 0.1.
In addition to the possibilities mentioned in Section VII, violation of the Type II bounds might also be attributed to the EFT operators $\dot{\pi}^{3}$ and $\dot{\pi}(\partial_{i}\pi)^{2}$, whose contributions we have neglected. This contrasts with the Type I bounds, which cannot be violated by the leading order inflaton self-interactions induced by the sign-definite operator $(\partial\phi)^{4}$.

### v2

@@ VI Examples of explicit scenarios
Our bounds must be satisfied by any bispectrum resulting from the tree-level exchange of a unitary sector coupled to the inflaton field as specified in (11). The simplest scenario to test involves a linear superposition of a finite number of resonances in the hidden sector. In Fig. 6, we illustrate a few examples with three such resonances and confirm that the resulting bispectrum shapes lie within the allowed region. We emphasize again that, as shown by the red curve in the figure, the bispectrum can indeed change sign for sufficiently squeezed triangle configurations. This behavior arises from the cosmological collider oscillations dominating over the local contributions to the bispectrum (see Eq. (31)).
@@
As another non-trivial example, we take the hidden sector to be a conformal field theory(CFT) defined in the bulk of de Sitter space, with $\cal O$} representing a scalar CFT operator of the scaling dimension $\delta$. (Such CFT should not be confused with a hypothetical CFT defined on the future boundary of de Sitter space, which might be dual to the theory in the bulk Strominger (2001).)
@@
$\times\Gamma(\delta-3/2-i\mu)\,\Gamma(\delta-3/2+i\mu)\,.$ (53)
@@
In the case of a CFT operator with conformal dimension $\delta$, the spectral density asymptotically behaves as $\rho(\mu)\propto\mu^{2\delta-3}$. This implies that, for the convergence of the spectral integral, we must have $\delta<2$. In conclusion, we expect our bounds to hold for the exchange of scalar operators with dimensions in the range $3/2<\delta<2$, where the two-point function can be expressed as a UV-finite integral over the mass spectrum of heavy fields in principal series.
@@
In Fig. 7, using the spectral density given in Eq. (53), we have numerically evaluated the bispectrum arising from the exchange of a scalar CFT operator for three sample values of the conformal dimension $\delta\in(3/2,2)$. As shown in the plots, the resulting shapes lie within the allowed region.
@@ VIII Typical non-Gaussianity templates and beyond
It should be emphasized that although the equilateral and orthogonal templates themselves are outside the allowed region, they still provide a useful basis for effectively capturing many of the curves within the peninsula (corresponding to different choices of the spectral density $\rho(\mu)$). As a result, existing observational bounds on $f_{\text{NL}}^{\text{eq}}$ and $f_{\text{NL}}^{\text{orth}}$ can already be used to constrain the amplitude of such curves, along the lines of Cabass et al. (2024).
Looking ahead, it would be interesting to identify a sufficient set of templates to effectively capture the diversity of possible shapes encompassed by the peninsula (in this direction, see the recent work Sohn et al. (2024) where a variety of templates associated with the exchange of massive fields are compared with the Planck CMB data—though some of the studied examples lie outside the scope of our assumptions, for instance due to strong breaking of boosts Pimentel and Wang (2022); Jazayeri et al. (2023); Jazayeri and Renaux-Petel (2022)). Beyond the standard template-based analyses, it is also important to consider how best to analyze the bispectrum data to assess whether it favors or disfavors the carved region as a whole. We leave this question for future investigations.
@@ IX Future Directions
Figure 8: The figure shows that the standard non-Gaussianity templates (Local, Orthogonal, and Equilateral) are inconsistent with our bounds, while the shape generated by the leading-order EFT operator $\delta\dot{\phi}(\partial\delta\phi)^{2}$ lies within our bispectrum island. See the discussions in Section VIII. Note that, compared to the logarithmic x-axis of Fig. 4, the horizontal axis is linear in $k_{3}/k_{1}$. The pink area represents the same region between the two curves in Fig. 4, which are multiplied by $(k_{3}/k_{1})^{-1/2}$ on the y-axis.
@@
Acknowledgments We would like to thank Giovanni Cabass, Sebastián Céspedes, Calvin Yi-Ren Chen, Carlo Contaldi, Enrico Pajer, Guilherme L. Pimentel, Arthur Poisson, David Stefanyszyn, Sébastien Renaux-Petel, Xi Tong, Kamran Salehi Vaziri, Dong-Gang Wang, Denis Werth, Chen Yang and Yuhang Zhu for stimulating discussions. We also thank Borna Salehian for comments on the manuscript. CdR and AJT are partially supported by the STFC Consolidated Grant ST/X000575/1. CdR and SJ are also supported by Simons Investigator Award 690508.
@@ Appendix A Large Boost Breaking Interactions
Figure 10: The left panel shows that in the Type II setup studied in Appendix A, the bispectrum template $B(k_{i};\mu)$ evaluated at the equilateral configuration changes sign as $\mu$ varies. In contrast, for certain triangle shapes—such as (1,1,0.1)(1,1,0.1)—the sign remains fixed for all values of $\mu$ (right). See Appendix A for further discussion.
@@
$M_{\mathrm{Pl}}\to\infty\,,\,\,\dot{H}\to 0\,\,\,(M_{\mathrm{Pl}}^{2}|\dot{H}|=\text{const.})\,$, (63)
@@
Notably, unlike the Type I setup, in Type II we cannot establish any definite sign for the bispectrum based on the positivity of $\rho(\mu)$. The reason is that the overall coupling $\kappa_{*}$ appearing above is not sign definite. In fact, even with a definite sign for $\kappa_{*}$, it is not possible to put bounds on the sign of the equilateral bispectrum in Type II, because the template $B(1,1,1;\mu)$ changes sign as $\mu$ is varied, see Fig. 10. This is also a roadblock to putting bounds on the bispectrum shape, conventionally normalized at the equilateral configuration. Such object would be unbounded for the same reason.
@@
which is expected to be bounded from above and below. Akin to the bispectrum shape in Type I, at a fixed momentum configuration, the shape $\tilde{{\cal S}}(k_{i})$ is also optimized by a spectral density that sharply picks at a particular mass index $\mu=\mu_{\text{max,min}}$, which we find numerically. In other words, we find the optimal values of the ansatz
@@
by varying the mass across the entire principal series, i.e. $\mu\in[0,\infty)$ (see the results in Fig. 11). As shown in Fig. 9, unlike the bounds obtained in the Type I setup, the maximum value of $\tilde{{\cal S}}$ is not set by the bispectrum associated with the exchange of the lightest field (i.e. $\mu=0$). Instead, for configurations with $k_{3}/k_{1}\gtrsim 0.1$, we find that $\mu_{\text{max}}\approx 1.9$ while $\mu_{\text{min}}=0$. For more squeezed triangles, however, $\mu_{\text{max}}$ decreases to zero, while $\mu_{\text{min}}$ varies from 0 in the squeezed limit up to $\approx 1.6$ as $k_{3}/k_{1}$ approaches 0.1.
In addition to the possibilities mentioned in Section VII, violation of the Type II bounds might also be attributed to the EFT operators $\dot{\pi}^{3}$ and $\dot{\pi}(\partial_{i}\pi)^{2}$, whose contributions we have neglected. This contrasts with the Type I bounds, which cannot be violated by the leading order inflaton self-interaction induced by the sign-definite operator $(\partial\phi)^{4}$.